OpenStudy (anonymous):

2x^2-2x-4=0 ? HELP ? D;

OpenStudy (anonymous):

can you draw it

OpenStudy (anonymous):

this question didn't come with a graph ? :o

OpenStudy (anonymous):

\[2x^2-2x-4 = 0\]

OpenStudy (anonymous):

2x^(2)-2x-4=0 Factor out the GCF of 2 from 2x^(2)-2x-4. 2(x^(2)-x-2)=0 Factor the trinomial x^(2)-x-2. 2(x+1)(x-2)=0 Divide both sides of the equation by 2. Dividing 0 by any non-zero number is 0. (x+1)(x-2)=0 Set each of the factors of the left-hand side of the equation equal to 0. x+1=0_x-2=0 Since 1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 1 from both sides. x=-1_x-2=0 Set each of the factors of the left-hand side of the equation equal to 0. x=-1_x-2=0 Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides. x=-1_x=2 The complete solution is the set of the individual solutions. x=-1,2

OpenStudy (anonymous):

Woah :O

OpenStudy (anonymous):

Do you understand @simba ? If not, I can walk you through it.